OpenStudy (anonymous):

1. What is the sum of the geometric sequence 8, –16, 32 … if there are 15 terms?

OpenStudy (anonymous):

how do i do this?

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

use the formula Sum = a r^n - 1 -------- r - 2 where a = first term, n = number of terms and r = common ratio common ration = -16 / 8 = -2

OpenStudy (anonymous):

how did u know the common ratio

OpenStudy (anonymous):

8-2^15-1 --------- -2-2 ?

OpenStudy (anonymous):

(8 * -2^15) - 1 --------- -2-2

OpenStudy (anonymous):

terms in a GP are of the form a,ar,ar^2,ar^3 so r = common ratio = ar /a in the above a = 8 and ar = -16

OpenStudy (anonymous):

sum =[ 8* -32768 - 1] / -4

OpenStudy (anonymous):

65536.25 \ - sum of GP with a common ratio integer >|1| gets bigger pretty quickly

OpenStudy (anonymous):

need any more help?